OpenStudy (anonymous):

Simultaneous equations. x+3y=6 x-y=2 No graphs. I solved Y by this: simplifying x+3y=6 X=6-3y inserting it into the other equation: 1(6-3y)+y=2 6-4y=2 6-2=4y y=1 But I need to find x. The answer is suppose to be (x,y)--> (3,1) Thank you.

OpenStudy (anonymous):

You made a mistake when you tried to find y. It should be: 1(6-3y)-y=2 6-4y=2 -4y=-4 y=1 substitute y=1 into x-y=2 you get: x-1=2 x=3 so (3,1)

OpenStudy (anonymous):

Ohh, I see. Then it makes sense. All because of the ''positive'' sign :-) Thank you, now I'll be able to solve them without any problem :-D